(x*x*x*x -81) / ( x -3)
Answer by terbear1963(43)

(x^4 - 81) / (x-3) = (x^2 - 9)(x^2 + 9) / x - 3 = (x-3)(x+3)(x^2+9)/x-3 = (x+3)(x^2+9) I HOPE THIS HELPS YOU OUT?! YOUR MATH FRIEND, TERRY cats7051@yahoo.com